I have 30 sweets and 15 chocolates in a basket,If I take five sweets out of the basket, what is the ratio (in it's simplest form) of the number of chocolates to the number of sweets left in the basket?

1
Expert's answer
2022-03-08T13:02:03-0500

"R=\frac{15}{30-5}=\frac{15}{25}=\frac{3}{5}."